sam reads 36 pages of a novel in 2 hrs at the same rate find the number of hrs it would take him to read the remaning 135 pages.
Answer by stanbon(75887)

sam reads 36 pages of a novel in 2 hrs at the same rate find the number of hrs it would take him to read the remaning 135 pages.
----
x/135 = 120/36
---
x = 135(120/36)
----
x = 450 minutes = 7 1/2 hours
=================================
Cheers,
Stan H.
==================